OpenStudy (anonymous):

If (f + g)(x) = 3x2 + 2x – 1 and g(x) = 2x – 2, what is f(x)? 3x2 + 1 3x2 – 1 3x2 + 4 3x2 – 4 x2 – 2

Directrix (directrix):

(f + g)(x) - g(x) = f(x)

Directrix (directrix):

3x^2 + 2x – 1 - (2x – 2) = ?

OpenStudy (anonymous):

i got 3x^2+4x-3

Directrix (directrix):

I did not get that.

Directrix (directrix):

You did not distribute the -1 in front of the parenthesis.

Directrix (directrix):

- (2x – 2) and 2x -2 are not the same thing.

OpenStudy (anonymous):

so its 3x^2+1 ?

Directrix (directrix):

That is correct.